Herschel Status Report - February 2011
The spacecraft continues to be in good health and is operating nominally. The issue of the autonomous shut down of the main star tracker (STR-1) twice in a row before Christmas (see previous status report), has been operationally resolved by disabling the check that led to the shut down.
On 21 February, in follow up to similar tests in October last year, a revised pointing test was conducted to improve the repeatability of pointings in nodded observations to sub-arcsecond level. The results from this test are being analysed.
Instrument operations have been nominal for the entire reporting period, apart from two events.
While routinely patching a HIFI memory section to correct a single-event-upset induced bit flip in the instrument's Local Oscillator Control Unit, the same memory segment seemed to be corrupted again as soon as it was patched. At the same time it was observed that the commands from the ground station were leaving in a burst-like fashion instead of the usual continuous stream of packets. It is suspected that due to instrument on-board software that can only cope with the arrival of data at a certain maximum rate, the chance occurrence of a bursty command stream during a patch that requires updating on-board memory is what led to the observed behaviour. The HIFI teams are working on confirming this suspected scenario in the laboratory. The MOC have refined the routine patch procedure to avoid the suspected scenario in the future, by interlocking commands released during the patch operation with a signal from the ground station that the previous packet has been dispatched to the spacecraft.
Some observing time was lost mid-February during a series of HIFI observations. Although the particular anomaly has not been seen before during the mission, the same effect has been seen in ground testing. It is a matter of instability at a transition between two frequency domains in the instrument which involve different, and discontinuous, tolerances for drifts in specific voltages. As a precautionary measure, all observations of the type that caused the anomaly (band 1a frequency switching) have been disabled for the time being until modified tolerances are validated by HIFI experts and uplinked to Herschel.
Ground Segment operations have been nominal and 100% of the data continue to be recovered. As of 23 February 2011, the completion of the different programme parts was:
|KPGT||Key Programme Guaranteed Time||:||69%|
|KPOT||Key Programme Open Time||:||66%|
|GT1||First in-flight Guaranteed Time||:||42%|
|OT1||First in-flight Open Time||:||0.6%|
For more details of these different programme parts, see the "overview of Herschel observing" linked from the right-hand menu.
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.
A routine station-keeping manoeuvre, or orbit correction manoeuvre, was performed on 11 February and had a nominal magnitude (delta-v) of 61.3cm/s. These routine manoeuvres ensure the spacecraft maintains its correct orbit about L2.
The ground segment is operating nominally. Data products are generated routinely and ingested into the Herschel Science Archive (HSA).